Help! I need the answer to this problem: A soccer field is 75 yards shorter than 3 times its width. Its perimeter is 370 yards. Find its dimensions.
Answer by vasu2qute(17)

let the width be x yards
length=3x-75
perimeter=2(length+width)=2(x+3x-75)=
2(4x-75) = 370
4x-75 =370/2
4x-75 =185
4x=185+75
4x=260
x=260/4=65
therefore width=65yards , length= 3x-75=3x65-75=195-75=120yards